Edit ModuleShow Tags
Edit ModuleShow Tags

Repsol Nabors Blowout at Qugruk 2 UPDATE Situation Report No. 5



Division of Spill Prevention and Response
Prevention and Emergency Response Program

Repsol Q2 Pad Gas and Mud Release

SPILL NUMBER: 12399904601

TIME/DATE OF SPILL: Repsol reported the incident to ADEC at 10:29 AM on February 15, 2012.

TIME/DATE OF SITUATION REPORT: 2:00 PM on February 18, 2012

TYPE/AMOUNT OF PRODUCT SPILLED: Repsol estimates Mud-water off south end of pad. Date: 2/17/2012 (PHOTO: ADEC-W. Ghormley)that approximately 42,000 gallons (1,000 barrels) of fresh-water-based drilling mud have been released to the ice drilling pad and adjacent snow-covered tundra. An unknown amount of gas has been released. The mud that was released is a mixture of minerals (primarily barium sulfate) and water. Drilling mud is used to lubricate the drill, control formation pressure, and remove cuttings from the bore. No oil has been spilled.

Mud-water off south end of pad on Feb. 17, 2012. (PHOTO: ADEC-W. Ghormley)

INCIDENT LOCATION: Qugruk #2 Pad, on the Colville River Delta, approximately 18 miles northeast of Nuiqsut and approximately 150 miles southeast of Barrow (70° 27’ 19” N, 150° 44’ 52” W).

CAUSE OF SPILL: Repsol contractor Nabors Drilling was drilling an exploratory well when the drill penetrated a shallow gas pocket at a depth of 2,523 feet, resulting in a gas kick. The gas kick drove drilling mud out of the well and through the gas diverter, which extends 75 feet from the drilling rig, onto the ice pad and adjacent snow-covered tundra. Additional mud was pumped into the well in an attempt to control it, but that mud was also forced out by the gas.


RESPONSE ACTION: The well control team began last night and continues today to thaw equipment on the drilling rig and to evaluate its condition. The team is not yet able to estimate when the rig will be ready to return to service. Air monitoring continues to indicate safe working conditions around the rig.

The start of well fluids cleanup awaits completion of the well control work. Photographs from yesterday’s fixed-wing overflight of the drilling pad show a deposit that may include drilling mud, water and cuttings from the well extending 100 feet or more from the edge of the pad.

Suspected misting on the SW side of pad. Date: 2/17/2012 (PHOTO: ADEC-W. Ghormley)An additional area may have been misted with the same materials. Another overflight was conducted this afternoon by helicopter to provide North Slope Borough personnel and an ADEC responder a more detailed look at the spill site.

Suspected misting on the SW side of pad, Feb. 17, 2010. (PHOTO: ADEC-W. Ghormley)

Two ADEC personnel remain stationed at the Palm Pad forward command post. One will conduct a document inspection at the Qugruk #2 Pad as soon as today. Two U.S. Fish and Wildlife Service personnel are scheduled to arrive at Palm Pad at 2:00 PM today to determine whether a collared polar bear may be denning in the vicinity of the spill site.

Yesterday, in an incident unrelated to the Qugruk #2 Pad gas release, residents of Nuiqsut reported smelling natural gas in the village. The source of the odor was a vent on a building where methanethiol (also called mercaptan), a chemical used to give a characteristic odor to natural gas in the village’s gas distribution system, is stored.

An Alaska Clean Seas team detected no natural gas during air monitoring conducted today in Nuiqsut at Repsol’s request. In the interest of safety, the North Slope Borough has hired a contractor to continue air monitoring in the village.

SOURCE CONTROL: The well ceased flowing at about 9:00 PM on February 16. Preparations for a well control operation are ongoing.

Mud-water on pad and tundra. Date: 2/17/2012 (PHOTO: ADEC-W. Ghormley)RESOURCES AFFECTED: Snow-covered tundra and ice pad. Drilling mud can affect tundra plants by changing soil pH and salinity. Brackish water produced by the well may also increase soil salinity.

Mud-water on pad and tundra, Feb. 17, 2012. (PHOTO: ADEC-W. Ghormley)

FUTURE PLANS AND RECOMMENDATIONS: Cleanup cannot safely begin until the well is under control. The well control team will give cleanup personnel approximately 24 hours notice in advance of completion of the well control operation.

Plans are in place for delineating and cleaning up the spill, and necessary personnel crews and equipment are on standby.

A reduced incident management team will continue to monitor the operation from Anchorage and two ADEC responders will remain at Palm Pad until cleanup operations are imminent, at which time additional staff will be recalled.

Incident Commander: Jeremy Michels, O’Brien’s Response Management
F.O.S.C.: Matt Carr, EPA
S.O.S.C.: Steve Russell, ADEC
L.O.S.C.: Gordon Brower, North Slope Borough
Field SOSC: Wes Ghormley, ADEC

Q2 Pad SE side mud-water.  Date: 2/17/2012 (PHOTO: ADEC-W. Ghormley)

WEATHER: Today: Scattered flurries, areas of fog; high temperature -7°F; east wind 15 mph; wind chill to -37°F.
Tonight: Scattered flurries, areas of fog; low -17°F; east wind 15 mph; wind chill to -37°F.
Sunday: Partly sunny with areas of blowing snow, scattered flurries and fog; high -8°F; east wind to 22 mph; wind chill to -38°F.

Q2 Pad SE side mud-water, Feb. 17, 2012. (PHOTO: ADEC-W. Ghormley)

TIME/DATE OF THE NEXT REPORT: 2:00 PM on February 19, 2012

FOR ADDITIONAL INFORMATION CONTACT: Ty Keltner, Public Information Officer, ADEC (907) 465-5009

Photographs and other spill information will be available for viewing at: http://www.state.ak.us/dec/spar/perp/index.htm

Maps Aerial photo/map showing incident location (PDF 333K)


Edit Module

Add your comment: